You've just joined the investment-banking firm of JP Morgan. They've offered you two different salary arrangements. The salary will be received monthly. You can have $75,000 per year for the next two years, or you can have $64,000 per year for the next two years, along with a $20,000 signing bonus today. If the interest rate is 10% compounded monthly, which do you prefer?

How can I solve this using a financial calculator? If not possible, how can I solve it?
Answer #1

option 1

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 6250*((1-(1+ 10/1200)^(-2*12))/(10/1200))
PV = 135442.84

option 2

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 5333.33333333333*((1-(1+ 10/1200)^(-2*12))/(10/1200))
PV = 115577.89

Total = 115577.89+20000 =135577.89

choose option 2